OpenStudy (anonymous):

Make 'n' the subject of the formula: s=(2a/3)sq.rt (b2-n2), where 0

OpenStudy (anonymous):

i don't really understand the point

OpenStudy (anonymous):

u want just solve the eqn

OpenStudy (anonymous):

just a correction its (b sq. - n sq.) and yes I just want to solve the equation. Thnx in advance :)

OpenStudy (anonymous):

i think something wrong with the queastion

OpenStudy (anonymous):

\[s=\frac{ 2a }{ 3 }\sqrt{b ^{2}-n ^{2}} \]

OpenStudy (anonymous):

s^2= (4a/9)b^2 -n^2 (9/4a)a^2=b^2 -n^2 s^2=(9/4a^3)-n^2=b^2

OpenStudy (anonymous):

so u now just take the cubic eqn

OpenStudy (anonymous):

I still can't figure it out :(

OpenStudy (anonymous):

sorry i would help u if i know